Hello,
I need help with this multiple choice question. Detailed working out would be greatly appreciated.
For the line y=x to be tangent to the curve with equation y=k/x-1
A. k=1
B. k>1
C. k>-1/4
D. k=-1/4
E. 4k+1<0
This is my working out so far:
Tangent: y=x
Curve: y=k/x-1
Appreciate your help.

Answer by ikleyn(52781)

.

Your formula   y = k/x-1   is   AMBIGOUS,   i.e. can be read in two different ways.


One way to read it is   y = (k/x) - 1.

Another way is   y = k/(x-1).


Use parentheses in your formulas when you post to this forum
to show which part is the numerator and which is the denominator.